OpenStudy (anonymous):

John is 5 feet tall and casts a 7.5-foot shadow. His older sister Jane casts a 9-foot shadow. How tall is John’s older sister, in feet?

OpenStudy (anonymous):

This is a ratio problem: \[\frac{ 5 }{ 7.5 } = \frac{ x }{ 9 }\]and now just solve for "x"

OpenStudy (anonymous):

I think the answer is 6.5, am I wrong?

OpenStudy (anonymous):

Yes, multiply both sides of the equation by 9 and you'll get your answer.

OpenStudy (anonymous):

oh, let me try.

OpenStudy (anonymous):

i am still lost :(

OpenStudy (anonymous):

np, I'll go a little further here for you. Look again at my equation. It becomes:\[\frac{ 9\times5 }{ 7.5 } = x\]that becomes\[\frac{ 45 }{ 7.5 } = x\]and now just simplify the left side.

OpenStudy (anonymous):

\[\frac{ 45 }{ 7.5 }\]means 45 divided by 7.5, if that helps

OpenStudy (anonymous):

WOW!!!!!! I totally understand now. Thank you so much . I have to right this down

OpenStudy (anonymous):

Cool! It's great when that old lightbulb goes on!

OpenStudy (anonymous):

sure is :)

OpenStudy (anonymous):

I like the way you explained it rather than my teacher